OpenStudy (anonymous):

integration of csc^2 x*sec^2 x dx

OpenStudy (anonymous):

please type by latex

OpenStudy (anonymous):

\[\int\limits_{}^{} \csc ^{2}xsec ^{2}x dx\]

OpenStudy (anonymous):

\[\int\limits {\frac{dx}{\cos^2 x \sin^2 x}}=\int\limits {\frac{dx}{\cos^2 x}} + \int\limits {\frac{dx}{\sin^2 x}}\] Try to have a next step :)

OpenStudy (anonymous):

you have \[1=\sin^2 x +\cos^2 x\] in numerator, so we can divide into a sum above.

OpenStudy (anonymous):

Notice that \[ \csc ^2(x) \sec ^2(x)=4 \csc ^2(2 x) \\ \] This comes from the fact \[ \sin(2 x) = 2 \sin(x) \cos(x) \] Notice that the derivative of cot(2x) is \[ -2 \csc ^2(2 x) \] You can now deduce your integral.